OpenStudy (anonymous):

y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (anonymous):

can you substitute the value of y from one equation in for y in the other equation?

OpenStudy (anonymous):

yes? i think so

OpenStudy (anonymous):

ok try it

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

y= y-1?

OpenStudy (anonymous):

it says y = 3x + 3 and y = x - 1, so if y is equal to both of those things can you set those things equal to each other?

OpenStudy (anonymous):

yes because they are equal.. i think so, math is my worst:(

OpenStudy (anonymous):

its fine, so set them equal to each other, what do you get?

OpenStudy (anonymous):

y=3y+3??

OpenStudy (anonymous):

ok lets try a simpler problem, then come back to this one

OpenStudy (anonymous):

if y = 10 and y = x then doesn't 10 = x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so you have y = 3x + 3 and y = x - 1 so what is 3x + 3 equal to 3x + 3 = ?

OpenStudy (anonymous):